40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6746     Entry time: Sat Jun 2 03:19:37 2012
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: Y green beat note found? - too small 

Summary:
  I tried to find Y arm green beat in order to do the mode scan.
  I found a beat peak(see attached picture), but the amplitude seems too small.
  It is may be because the alignment/mode matching of the green beams at the PSL table is so bad. Or, the peak I found might be a beat from junk light.

What I did:
  1. Aligned Y arm to the IR beam from MC.

  2. Re-aligned Y end green beam to the Y arm using steering mirrors on the Y end table.

  3. Re-aligned PSL green optics.

  # C1:GCV-GREEN_TRY is temporary connected to the DC output of the Y green beat PD.

  4. Temperature of the PSL laser was 31.48 deg C, so I set "T+" of the Y end laser to 34.47 deg C, according to Bryan's formula (elog #4439);

  Y_arm_Temp_set = 0.87326*T_PSL + 6.9825

  5. Scanned Y end laser temperature by C1:GCY-SLOW_SERVO2_OFFSET. Starting value was 29725 and I scanned from 27515 to 31805, by 10 or 100. Laser frequency changes ~ 6 MHz / 10 counts, so it means that I scanned ~ 2.5 GHz. During the scan, I toggled C1:AUX-GREEN_Y_Shutter to make sure the green beam resonates in TEM00 mode.

  # I made a revolutionary python script for toggling channels(/opt/rtcds/caltech/c1/scripts/general/toggler.py). I made it executable.

  6. Found a tiny beat note when C1:GCY-SLOW_SERVO2_OFFSET = 29815. I confirmed it is a beat signal by blocking each PSL and Y arm green beam into the beat PD. I left  C1:GCY-SLOW_SERVO2_OFFSET = 29815.

  7. I found that Bryan's formula;

Y_arm_Temp_meas = 0.95152*T_PSL + 3.8672
Y_arm_Temp_set = 0.87326*T_PSL + 6.9825

  was actually

Y_arm_Temp_set = 0.95152*T_PSL + 3.8672
Y_arm_Temp_meas = 0.87326*T_PSL + 6.9825

  according to his graph(elog #4439). So, I set  "T+" of the Y end laser to 33.82 deg C.

  8. This time, I scanned PSL laser temperature by C1:PSL-FSS_SLOWDC. I found a tiny beat note when C1:PSL-FSS_SLOWDC = 1.0995. C1:PSL-FSS_SLOWDC has 10 V range, so I scanned ~ 10 GHz, assuming the laser frequency changes 1 GHz/K and the temperature changes 1 K/V.

  9. Re-aligned PSL green optics so that the beam hits optics at their center, and checked that the poralization of the two green beams are the same.

  10. Checked that amplifier ZFL-100LN+ on the beat PD is working correctly. The power was supplied correctly (+15 V) and measured gain was ~ 25 dBm.

  11. Exchanged BNC cable which connects the beat PD to the spectrum analyzer. Previous one we used was too long and it had -15 dB loss(measured). I exchanged to shorter one which has -2 dB loss.

Beat note amplitude estimation:
  The amplitude of the beat note observed in the spectrum analyzer was ~ -54 dBm. According to the estimation below, it seems too small.

  The measured power of the two green beams are

  P_Y = 4 uW
  P_PSL = 90 uW

  So, the power of the beat signal should be

  P_beat ~ 2 sqrt(P_Y * P_PSL) = 37 uW

  Responsivity and transimpedance of the beat PD (Broadband PD, LIGO-T0900582) are 0.3 A/W and 2 kOhm. So, the power of the electrical signal is

  W = (P_beat * 0.3 A/W * 2 kOhm / sqrt(2))^2 / 50 Ohm = 5 uW

  5 uW is -23 dBm. We have +25 dB amplifier after the PD and the loss of the BNC cable is -2 dB. So, if the two beams interfere perfectly, the peak height of the beat signal should be ~ 0 dBm. The measured value -54 dBm seems too small. According to elog #5860, measured value by Kiwamu and Katrin was -36 dBm.

Current values:
  PSL laser temperature: 31.48 deg C (PSL HEPA 100%)
  Y end laser "T+": 33.821 deg C
  Y end laser "ADJ": 0
  C1:GCY-SLOW_SERVO2_OFFSET = 29815 (was 29725)

Attachment 1: CIMG1437.JPG  2.993 MB  Uploaded Sat Jun 2 04:30:56 2012  | Hide | Hide all
CIMG1437.JPG
ELOG V3.1.3-